A potentiometer wire 10m long and having 20 Ω  resistance is connected in series with a battery and a resistance R = 2160 Ω . A Daniel cell of emf 1.08 v is balanced across the P.D. of the resistance R. The emf of thermo couple is balanced across 300 cm of the wire. The emf of thermo couple is 
 
 
 
 

Grade:12th pass

1 Answers

Eshan
5 years ago
Current through the potentiometer wire=i=\dfrac{E}{20+2160}

Hence potetial difference across the potentiometer wire=(\dfrac{E}{20+2160})(20)
Therefore, emf of thermocouple=potential difference over 3m length of wire=\dfrac{3}{10}\times(\dfrac{E}{20+2160})(20)=3mV